OpenStudy (vijay):

P and Q each stand for whole numbers. P + Q = 1000 P is 150 greater than Q. Calculate the numbers P and Q ?

OpenStudy (zarkon):

\[P=Q+150\] substitute and solve

OpenStudy (vijay):

what is the meaning of this P is 150 greater than Q.

OpenStudy (anonymous):

P is bigger than Q by 150 P = Q + 150

OpenStudy (anonymous):

P=Q+150 P+Q=1000 Q+150+Q=1000 2Q+150 = 1000 2Q=850 Q=425 and P =575

OpenStudy (anonymous):

Q + 150 + Q = 1000 2Q + 150 = 1000 2Q = 850 Q = 425

OpenStudy (vijay):

thanks, i got it now